OpenStudy (anonymous):

help:/ Determine if the number is written in scientific notation. If not explain. 1. 4.8x100^7 No; it is not written as a number times a power of 10 yes; the number is written in scienfic notation no; the first factor is not a number between 1 and 10 2. 5.1x10^-8 yes; the number is written in scienfic notation no; the first factor is not a number between 1 and 10 no; it is not written as a number times a power of 10

jimthompson5910 (jim_thompson5910):

scientific notation follows the basic format b x 10^n where b is some number between 1 and 10 (NOT including 10 though) and n is some integer

jimthompson5910 (jim_thompson5910):

some numbers in scientific notation are: 2.3 x 10^7 8.1 x 10^(-5) 1.078 x 10^(12)

jimthompson5910 (jim_thompson5910):

so what can you tell me about 4.8 x 100^7

OpenStudy (anonymous):

i think 1 is b and 2 is c

jimthompson5910 (jim_thompson5910):

why 1 b

jimthompson5910 (jim_thompson5910):

why do you think 4.8 x 100^7 is in scientific notation?

jimthompson5910 (jim_thompson5910):

the base is 4.8 that's between 1 and 10 and it's not equal to 10 so that part works

jimthompson5910 (jim_thompson5910):

but.... the 100 *should* be 10 if you look back to the basic format I wrote out

OpenStudy (anonymous):

so its a

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

and 2 is what c?

jimthompson5910 (jim_thompson5910):

the reason why 10 is chosen over 100 is because you can move the decimal one spot at a time (when you change the exponent by 1) but with 100, you would have to move the decimal point twice per every change in exponent by 1

jimthompson5910 (jim_thompson5910):

so that's why 100 doesn't really work out

jimthompson5910 (jim_thompson5910):

5.1 is that between 1 and 10 (and not equal to 10)?

OpenStudy (anonymous):

wait is it a?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

compare the basic form b x 10^n with 5.1x10^-8 and you'll see that b = 5.1 and n = -8

jimthompson5910 (jim_thompson5910):

b = 5.1 is between 1 and 10 and not equal to 10 you have a power of 10 and the exponent is some integer

jimthompson5910 (jim_thompson5910):

so it all works out

OpenStudy (anonymous):

can u help me with more quuestions?

jimthompson5910 (jim_thompson5910):

sure, just a few more though

OpenStudy (anonymous):

What is the order of 1.5x10^3, 1.4x10^-1, 2x10^3, and 1.4x10^-2 from, least to greatest? 1.4x10^2, 1.4x10^-1, 1.5x10^3, 2x10^3 1.4x10^-1,1.4x10^2, 2x10^3, 1.5x10^3 2x10^3, 1.5x10^3, 1.4x10^-1, 1.4x10^-2 2x10^-3, 1.4x10^2, 1.5x10^3,1.4x10^-1 i thinki think it is A ?

OpenStudy (anonymous):

2. Simplify this expression(-4)^-6(-4)^7 4^13 -4 13 1 is it A?

jimthompson5910 (jim_thompson5910):

one at a time please

jimthompson5910 (jim_thompson5910):

it might help to expand out each number 1.4x10^2 = 140 (move the decimal point 2 places to the right) 1.4x10^-1 = 0.14 (move the decimal point 1 place to the left) 2x10^3 = 2000 (move the decimal point 3 places to the right) 1.4x10^-2 = 0.014 (move the decimal point 2 places to the left)

jimthompson5910 (jim_thompson5910):

so... the list of numbers: 1.5x10^3, 1.4x10^-1, 2x10^3, and 1.4x10^-2 is really the same as this list of numbers (in the same order) 140, 0.14, 2000, 0.014

jimthompson5910 (jim_thompson5910):

so it's your job to sort 1.5x10^3, 1.4x10^-1, 2x10^3, and 1.4x10^-2 but it might be easier to sort 140, 0.14, 2000, 0.014

jimthompson5910 (jim_thompson5910):

from least to greatest

jimthompson5910 (jim_thompson5910):

oh i missed one...sorry about that 1.5 x 10^3 = 1500

jimthompson5910 (jim_thompson5910):

let me redo, one sec

jimthompson5910 (jim_thompson5910):

1.5x10^3, 1.4x10^-1, 2x10^3, and 1.4x10^-2 1.5x10^3 = 1500 1.4x10^-1 = 0.14 2x10^3 = 2000 1.4x10^-2 = 0.014

jimthompson5910 (jim_thompson5910):

I misread the initial list of numbers, but it's fixed now

OpenStudy (anonymous):

so its 0.14, 0.014,140,1500,2000

jimthompson5910 (jim_thompson5910):

sorry, 140 isn't part of the list I misread and thought it was

jimthompson5910 (jim_thompson5910):

so kick that out

jimthompson5910 (jim_thompson5910):

0.014 is smaller than 0.14 so flip the two

jimthompson5910 (jim_thompson5910):

so here is the correct order 0.014, 0.14, 1500, 2000 now convert back to scientific notation

jimthompson5910 (jim_thompson5910):

to convert back, you just undo what we did above (so if we went 3 places to the right, we just go 3 places back to the left) 0.014 = 1.4 x 10^-2 0.14 = 1.4 x 10^-1 1500 = 1.5 x 10^3 2000 = 2 x 10^3

OpenStudy (anonymous):

0.14, 0.014,1500,200

OpenStudy (anonymous):

2000

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

btw this problem has a few typos in it (I think, not 100% sure though) for instance, in the original list, you refer to 1.4 x 10^-2 (the exponent is negative) but in the first answer choice, you have 1.4 x 10^2 (but now the exponent is positive)

OpenStudy (anonymous):

so it si a

jimthompson5910 (jim_thompson5910):

i think so, if choice A is really 1.4x10^-2, 1.4x10^-1, 1.5x10^3, 2x10^3 then yes it is

OpenStudy (anonymous):

. Simplify this expression(-4)^-6(-4)^7 4^13 -4 13 1 is it A?

jimthompson5910 (jim_thompson5910):

The bases are the same so you can add the exponents

jimthompson5910 (jim_thompson5910):

-6+7 = ???

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

so the new exponent is now 1

jimthompson5910 (jim_thompson5910):

therefore, we get (-4)^1 = ???

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

any number to the first power is that same number

jimthompson5910 (jim_thompson5910):

x^1 = x

OpenStudy (anonymous):

so its i -4

jimthompson5910 (jim_thompson5910):

yep -4